### v1

# Pion Valence Structure at Intermediate x in the Residual Field Model

Abstract
We calculate the valence parton distribution function of the pion, for the range of $0.1\leq x\leq 0.6$, based on the soft dynamics of the partons within the valence-residual field model. Within this model we assume that the soft partonic structure of the pion is defined by the valence current $q\bar{q}$ cluster embedded in the residual field of the pion. The parton distribution function is calculated using phenomenological wave functions for the $q\bar{q}$ cluster and residual field. Our result indicate that the peak position of x weighted valence parton distribution depends on the mass of the residual system and the parameters of the system are obtained by fitting the height and the peak position of this distribution to extracted PDFs. This comparison reveals that unlike for the case of nucleon PDF, very little residual mass is needed to describe existing pion PDFs at intermediate x.
I Introduction
As the lightest hadron and a pseudo-Goldstone boson of chiral symmetry breaking, the pion holds a unique position in hadronic and nuclear physics. Understanding its structure is essential for understanding Quantum Chromodynamics (QCD) nature of strong interaction of nucleons which in part is mediated through the pion exchange. Exploring pion structure at different light-front momentum domain of quarks and gluons will help also to trace the the transition from the partonic to the hadronic degrees of freedom in nuclear interactions at extreme conditions.
One quantity that allows to explore the dynamical structure of pion is its parton distribution function (PDF). PDFs measured over a wide range of x, $10<x<$ 1 can allow to address such questions as the dynamics of the valence distribution as well the role of the sea-quarks and gluons in the pion. It is interesting to explore the role of the valence, sea-quarks and gluons in pion especially from the point of view of their role in mediating interactions between nucleons in nuclei.
Much progress has been made in recent years in the study of pion parton distribution functions (PDFs)Barry et al. (2018a, 2021a). Recent analysis resulted in a high quality valence quark distribution in the pion as well as produced an unexpected result for PDFs at $x\to 1$. At $x\to 1$ if hard gluon exchange plays dominating role in generating PDFs, then within perturbative QCD one predicts an analytic form of pion PDFs to be $(1-x)^{2}$arrar and Jackson (1975); Berger and Brodsky (1979); Gunion et al. (1984). However the mentioned analysis resulted in a behavior of $(1-x)^{\beta}$, with $\beta\approx 1.01$.
Another aspect of the mentioned analysis is obtaining high quality PDF’s at intermediate $x$. For intermediate $x$ (0.1-0.6) the nonperturbative QCD dynamics define the structure of parton PDFs. One prominent feature of valence parton PDFs in this region is that the x weighted partonic distribution ($xf_{i}(x)$) shows prominent peak associated with the average momentum fraction carried by valence quarks. As it was shown in our recent workLeon and Sargsian (2020, 2022) modeling such distribution allows to ascertain the role of the quark-gluon system residual to the valence quarks in the hadron. In Ref.Leon and Sargsian (2020) We developed a model within residual mean field approximation which introduces Light-Front wave functions of valence quark cluster embedded in the residual field of quark-gluon system described by residual LF wave function. Within this model for the case of nucleon it was found that at the starting $Q^{2}_{0}$ when one expects partonic picture to be relevant, the $xf_{i}(x)$ for valence quark PDFs of nucleon peaks at $x_{p}\leq{1\over 4}$ not at ${1\over 3}$ as naively expected. By describing these peaking properties of valence PDFs of nucleons, we found that it requires the residual system to have a mass comparable to pion mass. As such the model indirectly favored the existence of the structure in the nucleon resembling a pion cloud.
In the current work, we develop a residual field model for the description of pion focusing on the domain of $0.1\leq x\leq 0.6$ and investigating the possible properties of the residual system in the pion.
The model divides the hadron into a valence core and a residual structure while utilizing flavor number conservation to treat the valence structure as effectively a system with a fixed number of fermions. The outline for this work is as follows: Sec. II provides a description of the theoretical framework being used and a review of the residual field model. In Sec. III we derive the PDFs based on the model and in Sec. V we compare the numerical estimates of our results with recently obtained phenomenological parameterization of pion PDFs in the range of $0.1\leq x\leq 0.6$. Our focus in this comparison is to evaluate the mass and the distribution of possible residual system in the pion. Sec. VI contains conclusions and outlook. Technical details for certain calculations and specific mathematical definitions are presented in the Appendix.
II Theoretical Framework
In light front (LF) quantization one probes quantum fields at fixed light-cone time, $x^{+}=t+z$, which provides several simplifications in describing high energy processes Brodsky et al. (1998). The LF approach provides a framework to study hadrons in terms of their light front wave functions (LFWFs). In principle, knowing the LFWFs would allow us to specify completely a hadronic state. However, finding the LFWFs would involve solving an infinite set of coupled integral equations, and thus often in practice they are modeled. The modeling of the valence LFWFs for the pion has been done to study the distribution amplitude of the pion Huang et al. (1994). There has also been modeling on the pion based on AdS/CFT correspondence and light front holography Trawiński (2016a, b).
A hadron in a ket state represented by $|P\rangle$ can be described through an expansion in terms of the LFWFs ($\psi_{i/N}\left(\{x_{i},\mathbf{k}_{i}\}_{i}\right)$) Brodsky et al. (1998); Kovchegov and Levin (2012):
$|P\rangle=\sum_{i}\int[d\mu_{i}]\psi_{i/N}\left(\{x_{j},\mathbf{k}_{j}\}_{j}\right)|i\rangle$ (1)
where the sum is over all Fock states $|i\rangle$ and $[d\mu_{i}]$ is the relevant integration measure. Color, spin, etc. degrees of freedom have been suppressed.
One can relate the experimentally measurable PDFs to the LFWFs to constrain them. In the $\overline{MS}$ factorization scheme, the PDF for a quark of flavor $i$ is defined as Sterman et al. (1995); Belitsky and Radyushkin (2005):
$f_{i}(x)=\int\frac{dz^{-}}{2\pi}e^{ixP^{+}z^{-}}\langle P|\bar{\psi}_{i}(z)\gamma^{+}\psi_{i}(0)|P\rangle|_{z^{+},z^{j}_{\perp}=0}$ (2)
Where we have employed the use of the light cone gauge $A^{+}=0$.
Our approach in this work is phenomenological, instead of taking an infinite Fock state expansion as in Eq. 1, we approximate the $|P\rangle$ state through the transition of $N\rightarrow VR$, where $V$ is a valence core and $R$ is a residual state and define phenomenological wave functions in such way that they represent the solution of Weinberg type equation on light-front. Such wave functions are inherently non-perturbative and need to be modeled based on the prominent features of the structure of the hadrons.
While quark number is not conserved, flavor is. The residual field model utilizes this fact by separating the hadron into a core of valence fermions whose number is conserved and a residual structure which contains everything besides the valence quarks, such as sea quarks and gluons and possibly a meson cloud in the case of the nucleon. This picture is expected to hold for moderate to high x, $x\gtrsim 0.1$. At lower $x$, Regge dynamics dominate and the model likely ceases to be valid. According to residual field model the deep inelastic scattering from a hadron (baryons or mesons) in leading approximation can be described by the Light-Front diagram of Fig.(1(a)) in which scattering takes place with the valence quark from the cluster. The consideration of the residual system in the scattering is equivalent to the spectral function representation of the partonic structure of hadron in which case valence quark distribution indirectly related to the dynamical characteristics of the residual system.
III The Pion in the Residual Field Model
III.1 General Formalism of Deep-Inelastic Scattering
We will consider probing the pion through deep inelastic electron scattering ($e\pi\rightarrow eX$) aimed at probing partonic distribution of valence quarks in the pion. Even though we consider a specific $\pi(e,e^{\prime})X$ reaction the PDFs evaluated in this case are universal and will be the same as PDF’s extracted from Drell-Yan processes in $\pi N\to l^{+}l^{-}+X$ reactions.
Within one-photon exchange approximation the scattering takes place with virtual photon $\gamma^{*}$, with four momenta $q^{\mu}(q_{0},\vec{q})$ where $q_{0}=E_{e}-E^{\prime}_{e}$ and $\vec{q}=\vec{k}-\vec{k}^{\prime}$ with $k^{\mu}(E_{e},\vec{k}_{e})$ and $k^{\mu\prime}(E^{\prime}_{e},\vec{k}^{\prime})$ being four momenta of initial and scattered
The cross section of DIS process in this can be expressed through the contraction of leptonic and hadronic tensors as follows:
${d\sigma\over dQ^{2}dx}={2\pi\alpha^{2}\over Q^{4}}{y^{2}\over Q^{2}}m_{\pi}L_{\mu\nu}W_{\pi}^{\mu\nu}$ (3)
where $Q^{2}=(q^{\mu}q_{\mu})$, $y={p_{\pi}q\over p_{\pi}k_{e}}$ and
$L_{\mu\nu}=4\left[\left(k_{\mu}-{q_{\mu}\over 2}\right)\left(k_{\nu}-{q_{\nu}\over 2}\right)\right]+Q^{2}\left[-g_{\mu\nu}-{q_{\mu}q_{\nu}\over Q2}\right],$ (4)
and the hadronic tensor in general can be defined from deep inelastic electromagnetic current $J^{\mu}(p_{X},s_{X},p_{\pi},s_{\pi})$ in the form:
$W_{\pi}^{\mu\nu}={1\over 4\pi m_{\pi}}\int\sum\limits_{X}\sum\limits_{s_{x}}(J^{\mu})^{\dagger}(p_{X},s_{X},p_{\pi},s_{\pi})J^{\nu}(p_{X},s_{X},p_{\pi},s_{\pi})(2\pi)^{4}\delta^{4}(q+p_{\pi}-p_{X})$
$\delta(p_{X}^{2}-M_{x}^{2}){d^{4}p_{x}\over(2\pi)^{3}}.$ (5)
Further, we will work with light-front momenta for which the 4-vector components are defined as follows:
$k^{\mu}=(k^{+},k^{-},\mathbf{k}^{\perp})=\left(k^{0}+k^{3},k^{0}-k^{3},(k^{1},k^{2})\right)$ (6)
where ‘3’ defines the direction of the large momentum component.
Furthermore we will perform calculations in the Drell-Yan-West reference frame where the 4-momenta of the pion, $p_{\pi}^{\mu}$ , and the virtual photon, $q^{\mu}$, are
$p_{\pi}^{\mu}=\left(p_{\pi}^{+},\frac{m_{\pi}^{2}}{p_{\pi}^{+}},\mathbf{0}^{\perp}\right),\ \ \ q^{\mu}=\left(0,\frac{2p_{\pi}\cdot q}{p_{\pi}^{+}},\mathbf{q}^{\perp}\right).$ (7)
where $Q^{2}=q_{\perp}^{2}$. The high momentum nature of the reaction is reflected in the fact that $p_{\pi}^{+}\gg\sqrt{Q^{2}}\gg m_{\pi}$.
With LF momenta defined in such way the $F_{2}$ structure function of pion is related to the +++++ + component of hadronic tensor as follows:
$F_{2}(x,Q^{2})={m_{\pi}Q^{2}\over 2x(p_{\pi}^{+})^{2}}W_{\pi}^{+,+}$ (8)
and within the leading approximation from the parton model
$F_{2}(x,Q^{2})=\sum\limits_{i}e_{i}^{2}xf_{i}(x,Q^{2}),$ (9)
were $i$ runs for both quarks and anti-quarks in the pion. In the case of the structure function it is dominated by valence quark contribution using the above relations allows to extract sum of quark and anti-quark valence quark distribution weighted by $x$. Usually invoking isospin symmetry it is assumed that the up valence distribution in $\pi^{+}$ is equal to the down valence distribution in $\pi^{-}$: $f_{V}(x,Q^{2})=u^{\pi^{+}}_{V}(x,Q^{2})=d^{\pi^{-}}_{V}(x,Q^{2})$.
III.2 Hadronic Tensor and The Scattering Amplitude
For calculation of pion structure function we consider the diagram of Fig.1(b) in which it is assumed that pion transitions to the cluster of valence $q\bar{q}$ state and residual $R$ state before the interaction of virtual photon with one of the valence quarks in the cluster.
(a) General Hadron
(b) Pion
Figure 1: Feynman Diagram for Deep Inelastic Scattering in Residual-Mean Field model
Assigning to the diagram of Fig. 1(b) an amplitude $A^{\mu}$, for the hadronic tensor, taking into account of the phase spaces of final state quarks and residual system one obtains:
The hadronic tensor, $W^{\mu\nu}$, has the form:
$W^{\mu\nu}(x,Q^{2})=\frac{1}{4\pi m_{\pi}}\sum_{\{h_{i},\tau_{i}\}}\int\delta(k_{R}^{2}-m_{R}^{2})\frac{d^{4}k_{R}}{(2\pi)^{3}}\delta(k^{\prime,2}_{1}-m^{2})\frac{d^{4}k^{\prime}_{1}}{(2\pi)^{3}}\delta(k_{2}^{2}-m^{2})\frac{d^{4}k_{2}}{(2\pi)^{3}}$
$(2\pi)^{4}\delta^{(4)}(p_{\pi}+q-\sum_{i=1}^{3}k_{i}-k_{R})A^{\mu\dagger}A^{\nu}$ (10)
where the sum $\sum_{\{h_{i},\tau_{i}\}}$ is over all the spins of the valence quarks and the residual system. The Lorentz invariant phase space measure in light front coordinates is,
$\delta(k^{2}-m^{2})d^{4}k$ $=$ $\frac{1}{2}\delta(k^{+}k^{-}-k_{\perp}^{2}-m^{2})dk^{-}dk^{+}d^{2}\mathbf{k}_{\perp}$ (11)
$=$ $\frac{dk^{+}d^{2}\mathbf{k}_{\perp}}{2k^{+}}|_{k^{-}=\frac{k_{\perp}^{2}+m^{2}}{k^{+}}}=\frac{dxd^{2}\mathbf{k}_{\perp}}{2x}|_{k^{-}=\frac{k_{\perp}^{2}+m^{2}}{xP_{\pi}^{+}}}$
where $x=\frac{k^{+}}{P_{\pi}^{+}}$ is the light front momentum fraction of pion constituent with momentum $k^{+}$. For energy-momentum conservation one obtains:
$\delta^{(4)}(P_{\pi}+q-k^{\prime}_{1}-k_{2}-k_{R})=$ $2\delta(p_{\pi}^{+}+q^{+}-k^{\prime,+}_{1}-k^{+}_{2}-k_{R}^{+})\delta(p_{\pi}^{-}+q^{-}-k^{\prime,-}_{1}-k^{-}_{2}--k_{R}^{-})$ (12)
×$\times\delta^{(2)}(\mathbf{p_{\pi}^{\perp}}+\mathbf{q}^{\perp}-\mathbf{k^{\prime\perp}_{1}}-\mathbf{k_{2}^{\perp}}-\mathbf{k}_{R}^{\perp}).$
Taking into account of Eq.(7) according to which $q^{+}=0$ and $k^{\prime+}_{1}=k_{1}^{+}$ one obtains:
$\delta(p_{\pi}^{+}+q^{+}-k^{\prime+}_{1}-k^{+}_{2}-k_{R}^{+})=\frac{1}{p_{\pi}^{+}}\delta(1-x_{1}-x_{2}-x_{R}),$ (13)
where $x_{1}={k_{1}^{+}\over p_{\pi}^{+}}$, $x_{2}={k_{2}^{+}\over p_{\pi}^{+}}$, and $x_{R}={k_{R}^{+}\over p_{\pi}^{+}}$. Using $k_{1}^{\prime}=p_{\pi}-k_{2}-k_{3}-k_{R}$ and neglecting initial transverse momenta compared to $q_{\perp}$,i.e. $k_{1,\perp}^{\prime 2}=(\mathbf{k}_{1,\perp}+\mathbf{q}_{\perp})^{2}\sim\mathbf{q}_{\perp}^{2}\gg k_{1,\perp}^{2}$ as well as using the condition $p_{\pi}^{+}\gg m_{\pi},m,\ m_{R},\ k_{1,2}^{\perp},\ k_{R}$ one evaluates:
$\delta(p_{\pi}^{-}+q^{-}-k^{\prime-}_{1}-k_{2}^{-}-k_{R}^{-})$ $=$ $\delta\left(\frac{m_{\pi}^{2}}{p_{\pi}^{+}}+\frac{2p_{\pi}\cdot q}{p_{\pi}^{+}}-\frac{k_{1,\perp}^{\prime 2}+m^{2}}{x_{1}p_{\pi}^{+}}-\frac{k_{2,\perp}^{2}+m^{2}}{x_{2}p_{\pi}^{+}}-\frac{k_{R,\perp}^{2}+m_{R}^{2}}{x_{R}p_{\pi}^{+}}\right)$ (14)
$\approx$ $\delta\left(\frac{2p_{\pi}\cdot q}{p_{\pi}^{+}}-\frac{Q^{2}}{x_{1}p_{\pi}+}\right)=\frac{x_{1}p_{\pi}^{+}}{2p_{\pi}\cdot q}\delta(x_{1}-x_{B}),$
where Bjorken x is defined as $x_{B}=\frac{Q^{2}}{2P\cdot q}$. Finally with the same conditions:
$\delta(p_{\pi}^{+}+q^{+}-k^{\prime+}_{1}-k^{+}_{2}-k_{R}^{+})=\delta(k^{+}_{1}+k^{+}_{2}+k_{R}^{+})$ (15)
Using Eqs.(13,14,15) in Eq.(12) for hadronic tensor in Eq.(10) one arrives at:
$W^{\mu\nu}(x,Q^{2})=\frac{1}{4m_{\pi}}\sum_{\{h_{i},\tau_{i}\}}\int[dx][d^{2}\mathbf{k}_{\perp}]{x_{1}\over p_{\pi}q}\delta(x_{1}-x_{B})A^{\mu\dagger}A^{\nu}$ (16)
where
$[dx]$ $=$ $\delta(1-x_{1}-x_{2}-x_{R})\frac{dx_{R}}{x_{R}}\prod_{i=1}^{2}\frac{dx_{i}}{x_{i}}$
$\left[d^{2}\mathbf{k}_{\perp}\right]$ $=$ $16\pi^{3}\delta^{(2)}(\sum_{i=1}^{2}\mathbf{k}_{i,\perp}+\mathbf{k}_{R,\perp})\frac{d^{2}\mathbf{k}_{R,\perp}}{16\pi^{3}}\prod_{i=1}^{2}\frac{d^{2}\mathbf{k}_{i,\perp}}{16\pi^{3}}$ (17)
As it follows from Eq.(8) and Eq.(9) to calculate the pion PDF one needs to evaluate the ”+” component of the scattering amplitude $A^{\mu}$.
III.3 Scattering Amplitude in Residual Field Model
To calculate the amplitude presented in Fig.(1(b)) we apply diagrammatic rules of Light-Front perturbation theory. Since momentum transferred to the quark/anti-quark in the pion is significantly larger than their initial momenta, the interference terms will be small and the cross section will represent the incoherent sum of the square of amplitudes corresponding to the scattering from quark and anti-quark. Thus we first calculate amplitude of scattering from quark and then use the result to evaluate the scattering from the anti-quark.
To calculate the amplitude we noticed that the diagram of Fig.(1(b)) is characterized by two intermediate states (indicated by vertical dashed lines) that are defined by two light-front propagators. With the momenta assigned for each particles in the scattering defined in the diagram one obtains:
$A^{\mu}(k_{1}^{\prime},h_{1}^{\prime},k_{2},h_{2},k_{R},h_{R})=\sum\limits_{h_{V}}{\bar{u}(k_{1}^{\prime},h_{1}^{\prime})ie_{q}\gamma^{\mu}(/\hskip-5.0ptk_{1}+m_{q})v(k_{2},h_{2})\Gamma^{V\to q\bar{q}}\chi_{V}\over k_{1}^{+}D_{2}}{\chi_{V}^{\dagger}\chi_{R}^{\dagger}\Gamma^{\pi\to VR}\chi_{\pi}\over k_{V}^{+}D_{1}},$ (18)
where $\Gamma^{\pi\rightarrow VR}$ is the effective vertex for the transition of pion into valence $q\bar{q}$ cluster and $\Gamma^{V\rightarrow q\bar{q}}$ describes the transition f the cluster to valence quark and anti-quark. Also, $h_{i}$ is the helicity of the i’th valence parton and $h_{1}^{\prime}$ is the helicity of scattered quark.
For the denominators $D_{1}$ and $D_{2}$ according to LF diagrammatic rules, one has:
$D_{1}=p_{\pi}^{-}-k_{R}^{-}-k_{V}^{-}={1\over p_{\pi}^{+}}\left(m_{\pi}^{2}-{k_{R\perp}^{2}+m_{R}^{2}\over x_{R}}-{k_{V\perp}^{2}+m_{V}^{2}\over x_{V}}\right)$ (19)
$D_{2}=p_{V}^{-}-k_{1}^{-}-k_{2}^{-}-={1\over K_{V}^{+}}\left(m_{V}^{2}-{k_{1\perp}^{2}+m^{2}\over\beta_{1}}-{k_{2\perp}^{2}+m^{2}\over\beta_{2}}\right).$ (20)
where $\beta_{i}=\frac{x_{i}}{x_{V}}$ for $i=1,2$ quark and anti-quark.
Furthermore, we use LF sum rule for the numerator of the propagator of off-energy shell quark with momentum $k_{1}$:
$(/\hskip-5.0ptk_{1}+m_{q})=\sum\limits_{h1}u(k_{1},h_{1})\bar{u}(k_{1},h_{1})+{\gamma^{+}(k^{2}-m^{2})\over k^{+}}$ (21)
and notice that since we are interested in the $A^{+}$ component of the scattering amplitude, but due to the relation $\gamma^{+}\gamma^{+}=0$, the second part of Eq.(21) does not contribute to the $+$ component of the amplitude. Inserting the first part of Eq.(21) into Eq.(18) and introducing phenomenological LF wave functions for valence $q\bar{q}$ cluster:
$\psi_{VR}(x_{V},\mathbf{k}_{R,\perp},x_{R},\mathbf{k}_{V,\perp})=\frac{\bar{\chi}_{V}\bar{\chi}_{R}\Gamma^{\pi\rightarrow VR}}{m^{2}_{\pi}-\frac{k^{2}_{V,\perp}+m^{2}_{V}}{x_{V}}-\frac{k^{2}_{R,\perp}+m^{2}_{R}}{x_{R}}}$ (22)
and residual system
$\psi_{q\bar{q}}(\beta_{1},\beta_{2},\mathbf{k}_{1,\perp},\mathbf{k}_{2,\perp})=\frac{\bar{u}(k_{1},h_{1})\Gamma^{V\rightarrow q\bar{q}}v(k_{2},h_{2})\chi_{V}}{m^{2}_{V}-\sum_{i=1}^{2}\frac{k^{2}_{i,\perp}+m^{2}_{i}}{\beta_{i}}}$ (23)
For the ”+”-component of the scattering amplitude assuming that the residual state has zero helicity and writing explicitly the total zero helicity state of the valence $q\bar{q}$ state as $\xi_{q\bar{q}}={u(k_{1},h_{1}){\bar{u}}(k_{2},-h_{1})-u(k_{1},-h_{1})\bar{u}(k_{2},h_{1})\over\sqrt{2}}$ one obtains:
$A^{+}(k_{1}^{\prime},h_{1}^{\prime},k_{2},-h_{1},k_{R})=\sum\limits_{h_{1}}\left[\bar{u}(k_{1}^{\prime},h_{1}^{\prime})ie_{q}\gamma^{+}u(k_{1},h_{1}){\psi_{q\bar{q}}(\beta_{1},\beta_{2},\mathbf{k}_{1,\perp},\mathbf{k}_{2,\perp})\over\sqrt{2}\beta_{1}}\right.$
$\left.-\bar{v}(k_{1},h_{1})ie_{\bar{q}}\gamma^{+}v(k_{1}^{\prime},h_{1}^{\prime}){\psi_{q\bar{q}}(\beta_{1},\beta_{2},\mathbf{k}_{1,\perp},\mathbf{k}_{2,\perp})\over\sqrt{2}\beta_{1}}\right]{\psi_{VR}(x_{V},\mathbf{k}_{R,\perp},x_{R},\mathbf{k}_{V,\perp})\over x_{V}}.$ (24)
By neglecting the masses of valence quarks, one evaluates only helicity conserving term of $\bar{u}(k_{1}^{\prime},h_{1}^{\prime})\gamma^{+}u(k_{1},h_{1})=\bar{v}(k_{1},h_{1})\gamma^{+}v(k_{1}^{\prime},h_{1}^{\prime})=2k_{1}^{+}\delta^{h^{\prime}_{1},h_{1}}$. This results in:
$A^{+}(k_{1}^{\prime},h_{1},k_{2},-h_{1},k_{R})=\left[2k_{1}^{+}ie_{q}{\psi_{q\bar{q}}(\beta_{1},\beta_{2},\mathbf{k}_{1,\perp},\mathbf{k}_{2,\perp})\over\sqrt{2}\beta_{1}}-2k_{1}^{+}ie_{\bar{q}}{\psi_{q\bar{q}}(\beta_{1},\beta_{2},\mathbf{k}_{1,\perp},\mathbf{k}_{2,\perp})\over\sqrt{2}\beta_{1}}\right]$
×$\times{\psi_{VR}(x_{V},\mathbf{k}_{R,\perp},x_{R},\mathbf{k}_{V,\perp})\over x_{V}}$ (25)
III.4 Valence Parton Distribution
Inserting Eq.(25) into the expression of hadronic tensor in Eq.(16) and using relations of Eq.(8) and (9) for the valence quark distribution one obtains:
$f_{V}(x_{B},Q^{2})$ $=$ $\int^{Q^{2}}\frac{dx_{1}d^{2}\mathbf{k}_{1,\perp}}{16\pi^{3}x_{1}}\frac{dx_{2}d^{2}\mathbf{k}_{2,\perp}}{16\pi^{3}x_{2}}\frac{dx_{R}d^{2}\mathbf{k}_{R,\perp}}{16\pi^{3}x_{R}}\delta(x_{1}-x_{B})16\pi^{3}\delta\left(1-x_{1}-x_{2}-x_{R}\right)$ (26)
$\times$ $\delta^{(2)}\left(\sum_{i=1,2,R}\mathbf{k}_{i,\perp}\right)\mid\Psi_{q\bar{q}}(\beta_{1},k_{1,\perp},\beta_{2},k_{2,\perp})\mid^{2}\mid\Psi_{VR}(x_{R},k_{R,\perp})\mid^{2}$
Expecting that Light-Front wave function of the $q\bar{q}$ system will depend on relative transverse momentum and relative light-front momentum fractions only, we can factorize above integrals by introducing transfer momentum of quark and anti-quark in the center of mass of the $q\bar{q}$ system
[…]
Substituting Eq.(27) into Eq.(26) one can factorize the integration by $k_{R,\perp}$. Furthermore integrating over $x_{2}$ results in
$f_{V}(x_{B},Q^{2})=\int_{0}^{1-x_{B}}\frac{dx_{R}}{(16\pi^{3})^{2}x_{B}(1-x_{B}-x_{R})x_{R}}$
×$\ \ \ \ \ \ \times\left[\int^{Q^{2}}d^{2}\tilde{\mathbf{k}}_{1,\perp}d^{2}\tilde{\mathbf{k}}_{2,\perp}\delta^{2}(\tilde{\mathbf{k}}_{1,\perp}+\tilde{\mathbf{k}}_{2,\perp})|\Psi_{q\bar{q}}(\beta_{1},\tilde{\mathbf{k}}_{1,\perp})|^{2}\int^{Q^{2}}d^{2}\mathbf{k}_{R,\perp}|\psi_{VR}(x_{R},\mathbf{k}_{R,\perp})|^{2}\right]$
Integrating now by $d^{2}\tilde{k}_{2,\perp}$ using the $\delta^{2}()$ function and renaming $\beta_{1}\equiv\beta$ and $k_{1,\perp}\equiv k_{\perp}$ one obtains:
$f_{V}(x_{B},Q^{2})=\int_{0}^{1-x_{B}}\frac{dx_{R}}{(16\pi^{3})^{2}x_{B}(1-x_{B}-x_{R})x_{R}}$
×$\ \ \ \ \ \ \times\left[\int^{Q^{2}}d^{2}\mathbf{k}_{\perp}|\Psi_{q\bar{q}}(\beta,{\mathbf{k}}_{\perp})|^{2}\right]_{q\bar{q}}\left[\int^{Q^{2}}d^{2}\mathbf{k}_{R,\perp}|\psi_{VR}(x_{R},\mathbf{k}_{R,\perp})|^{2}\right]_{VR}$
where $[\cdots]_{q\bar{q}}$ and $[\cdots]_{VR}$ we can interpret as response functions of valence - cluster and residual systems respectively.
IV Modeling Nonperturbative Light-Front Wave Functions
IV.1 General approach for two-body relativistic LFWF
To model $q\bar{q}$ cluster and $V-R$ wave functions we first note that the relativistic normalization of such a wave function is:
$\int\mid\Psi^{LF}(k_{12})\mid^{2}{dy\over y(1-y)}{d^{2}k_{12,\perp}\over 16\pi^{3}}=1$ (30)
where
$y={k_{1}^{+}\over p_{T}^{+}}$ (31)
with $k_{1}$ and $p_{T}$ being four momenta of one of the bound particles the target particle. Here the relative momentum in the CM of two particles are defined as
$k_{12}^{2}=\frac{\left(s_{12}-(m_{1}-m_{2})^{2}\right)\left(s_{12}-(m_{1}+m_{2})^{2}\right)}{4s_{12}}$
$={m_{1}^{2}(1-y)+m_{2}^{2}y+k^{2}_{12,\perp}\over 4y(1-y)}-{m_{1}^{2}+m_{2}^{2}\over 2}+{(m_{1}^{2}-m_{2}^{2})^{2}y(1-y)\over 4(m_{1}^{2}(1-y)+m_{2}^{2}y+k^{2}_{12\perp})}$ (32)
and
$s_{12}={m_{1}^{2}(1-y)+m_{2}^{2}y+k^{2}_{12,\perp}\over y(1-y)}$ (33)
We model LF wave function similar to Frankfurt and Strikman (1981); Sargsian and Vera (2023) assuming factorization of the relative momentum in the CM system it can be related to the non-relativistic wave function as follows:
$\Psi_{12}^{LF}(k_{12})=\Psi_{12}^{NR}(k_{12})\sqrt{{E_{1}E_{2}\over E_{1}+E_{2}}16\pi^{3}}$ (34)
where non-relativistic wave function is normalized as follows:
$\int\mid\Psi^{NR}(k_{12})\mid^{2}d^{3}k_{12}=1$ (35)
IV.1.1 Soft $q\bar{q}$ Wave Function
Assuming quark masses are the same, the above approach can be used to model the soft $q\bar{q}$ wave function in the form:
$\Psi^{LF}_{q\bar{q}}(\beta,k_{12,\perp})=\sqrt{8\pi^{3}E_{k}}\Psi^{NR}_{q\bar{q}}(k_{12}),$ (36)
where, $E_{k}=\sqrt{m^{2}+k_{12}^{2}}$ and $\vec{k}_{12}=(k_{12,z},\vec{k}_{12,\perp})$ is the relative momentum in the $q\bar{q}$ CM system, with the magnitude
$k_{12}=\sqrt{{m^{2}+k^{2}_{12,\perp}\over 4\beta(1-\beta)}-m^{2}}.$ (37)
For non-relativistic wave function we use two approximations, one Harmonic Oscillator type wave function
$\Psi^{NR}_{q\bar{q}}(k)=A_{V}\exp{\left[-\frac{B_{V}}{2}k_{12}^{2}\right]}$ (38)
and the other the ground state wave function of positronium.
A number of previous works have used the similar wave functions or a form similar to it Brodsky et al. (1981); Lepage and Brodsky (1980); Gunion et al. (1984); Trawiński (2016a); Pasquini and Rodini (2019) Using the above light-front wave function Eq. 36 for the $q\bar{q}$ system’s response function from Eq.29 one obtains:
$\left[\int^{Q^{2}}d^{2}\mathbf{k}_{\perp}|\Psi_{q\bar{q}}(\beta,{\mathbf{k}}_{\perp})|^{2}\right]_{q\bar{q}}=8\pi^{3}\int^{Q^{2}}d^{2}k_{\perp}E_{k}A_{V}^{2}e^{-B_{V}{k_{\perp}^{2}\over 4\beta(1-\beta)}}$
$=8\pi^{3}\int^{Q^{2}}d^{2}\mathbf{k}_{\perp}\sqrt{{k_{\perp}^{2}\over 4\beta(1-\beta)}}e^{-B_{V}{k_{\perp}^{2}\over 4\beta(1-\beta)}}={16\pi^{4}\sqrt{\pi}A_{V}^{2}\over B_{V}^{3\over 2}}\beta(1-\beta),$ (39)
where in the last step we neglected by the mass of the valence quarks and and assumed $Q^{2}\gg k_{\perp}^{2}$ (consistent with collinear approximation). In calculation of the above integral we set $Q^{2}\rightarrow\infty$.
One can see from the above equation that $q\bar{q}$’s response function peaks at $\beta={1\over 2}$
IV.1.2 Light-Front Wave Function of Valence-Residual System
To calculate the response function of residual system one has to model the the light-front wave function of valence-residual system. Here we use a Harmonic Oscillator (HO) type wavefunction in Eq.(34) obtaining
$\Psi_{12}^{LF}(x_{R},k_{R,\perp})=A_{R}\sqrt{16\pi^{3}}\sqrt{{E_{V}E_{R}\over E_{V}+E_{R}}}e^{-{B_{R}\over 2}k_{R}^{2}},$ (40)
where $E_{V}=\sqrt{m_{V}^{2}+k_{R}^{2}}$, $E_{R}=\sqrt{m_{R}^{2}+k_{R}^{2}}$. The $k_{R}$ is the 3-dimensional momentum defined in the center of mass of the $V-R$ system:
$k^{2}_{VR}=\frac{(s_{VR}-(m_{V}-m_{R})^{2})(s_{VR}-(m_{V}+m_{R})^{2})}{4s_{VR}},$ (41)
where
$s_{VR}=\frac{k_{R,\perp}^{2}+m_{R}^{2}}{x_{R}}+\frac{k_{R,\perp}^{2}+m_{V}^{2}}{1-x_{R}},$ (42)
resulting in
$k_{R}^{2}=$ ${m_{R}^{2}(1-x_{R})+m_{V}^{2}x_{R}+k^{2}_{R,\perp}\over 4x_{R}(1-x_{R})}$ (43)
$-{m_{R}^{2}+m_{V}^{2}\over 2}+{(m_{R}^{2}-m_{V}^{2})^{2}x_{R}(1-x_{R})\over 4(m_{R}^{2}(1-x_{R})+m_{V}^{2}x_{R}+k^{2}_{R\perp})}.$
Using above expressions for the $V-R$ wave function and relative momentum $k_{R}$, for the response function of the $VR$ system one obtains:
$R(x_{R})=\left[\int^{Q^{2}}d^{2}\mathbf{k}_{R,\perp}|\psi_{VR}(x_{R},\mathbf{k}_{R,\perp})|^{2}\right]_{VR}={16\pi^{3}}A_{R}^{2}\int^{Q^{2}}d^{2}\mathbf{k}_{R,\perp}{E_{V}E_{R}\over(E_{V}+E_{R})}e^{-B_{R}k_{R}^{2}}.$ (44)
Note that in the above equation masses of the valence cluster $m_{V}$, and and residual state $m_{R}$, are virtual and characterize the transition into non-stationary virtual state of a valence cluster and the residual system, thus they do not necessarily need to correspond to the pion mass. Additionally there is no condition for collinear approximation so the response function may exhibit some $Q^{2}$ dependence.
IV.2 Modeled Soft Pion PDF
Inserting Eq.(39) and Eq.44) into Eq.(29) for the modeled parton distribution of pion one obtains:
$f_{V}(x_{B},Q^{2})={(\pi)^{3\over 2}A_{V}^{2}A_{R}^{2}\over B_{V}^{3\over 2}}\int_{0}^{1-x_{B}}\frac{dx_{R}}{x_{R}}{1\over(1-x_{R})^{2}}\ \left[\int^{Q^{2}}d^{2}\mathbf{k}_{R,\perp}{E_{V}E_{R}\over(E_{V}+E_{R})}e^{-B_{R}k_{R}^{2}}\right]$
V Fitting Results
V.1 Parameters of Fitting
We use recent results from JAM18 collaborationBarry et al. (2018b, 2021b) on pion PDF to fit Eq.(45). Our main focus in the fitting procedure is to describe the position and the hight of the x weighted parton distribution in accordance with our model assumption that these features of valence PDFs are characterized by the soft partonic dynamics. As it follows from Eq.(45) the masses of residual, $m_{R}$, and valence, $m_{V}$, systems, as well as the width of the $V-R$ distribution define the position and the height of the peak.
We choose starting $Q^{2}=Q_{0}^{2}=1.69$ GeV2, where $Q_{0}$ is chosen in JAM18 parameterization to be the $c$-quark mass. The extensive fitting analysis indicates that the position of the peak can be reproduced only if the virtual mass of the residual system is close to zero. This is very different from our analysis of nucleon PDF’s Leon and Sargsian (2020); Leon et al. (2020) in which we found that the residual mass for the best fit is close to the pion mass, $m_{R}^{N}\approx 0.14$, indicating the validity of meson cloud picture of valence parton distribution in the nucleon.
The best fit of the JAM18 pion PDF is presented in Fig.2, for the parameters of $m_{R}=0$, $m_{V}=0.95$ GeV and $B_{V}=1.2$GeV-2. Note that the qualitative agreement of the data requires $m_{V}\sim Q_{0}$ and $m_{R}\sim 0$. As the figure shows, the soft mechanism describes the empirical PDF reasonably well in the range of $0.850.1<x<$ 0.85, leaving very little room for the hard mechanism for the generation of PDF at large x.
Figure 2: Description of x-weighted pion PDF at $Q_{0}=1.3$ GeV for $m_{R}=0$, $m_{V}=0.95$ GeV and $B_{R}=1.2$ GeV-2.
V.2 Interpretation of the Results
The parameters of the model that provide the best fit to the empirical PDF allows to accertain the mechanism dominating parton dynamics at the range where the peak of x-weighted PDF distribution is observed. For this we first analyze the response function of the residual system presented in Eq.(44). As Fig.3 shows the response function has a wide distribution at $x_{R}$ slightly peaking at $x_{R}\approx 0.5$.
Figure 3: The $x_{R}$ dependence of response function $R(x_{R})$ at different $Q^{2}$.
This indicates that the valence $q\bar{q}$ system carries a total momentum fraction $1-x_{R}\approx 0.5$. Since $x$-Bjorken corresponds to the Light-Front momentum fraction of interacting quark, then one concludes that the observed distribution indicates the dominance of the Feynman mechanismFeynman (1972) according to which the interacting quark carries the most of the Light-front momentum fraction with remaining momenta distributed among wee partons.
Another observation, as it was noted above, the result of $m_{R}\approx 0$, this indicates that the residual system does not exhibit any structure (as it was case of nucleon PDFs). On the other hand the large $m_{V}\approx 0.95$ GeV indicate the large virtuality of interacting quark comparable with the $Q_{0}$ of the probe. This together with the relatively small $B_{R}$ indicates that while interacting valence quark is highly virtual, the residual system is spread over the bulk size of the pion.
This picture is consistent with the phenomenology that nucleon interactions in a nucleus proceed through the pion exchange in which no quark-interchanges are involved.
Finally, the surprising result of the fit is that the considered model leaves little room for hard mechanism of high x PDF generated though the hard gluon exchange between valence $q\bar{q}$ system.
VI Conclusion and Outlook
We have calculated pion valence parton distribution within residual field model, which was previously applied to the calculation of Nucleon PDFs. The models assume no hard interaction between valence quarks and it separates $q\bar{q}$ valence cluster from the residual system. Both, $q\bar{q}$ and $V-R$ systems are described by nonperturbative Light-Front wave functions that satisfies relativistic normalization. These wavefunctions allow to introduce the residual response function as well as the response function of the $q\bar{q}$ valence cluster. The pion PDF represents the convolution of these response functions. We analyzed the parameters of the model by fitting the position and the height of the x-weighted PDF using empirical JAM18 distribution.
The fitting indicates that one can describe the above attributes of pion PDF only if the mass of the residual system is close to zero. However we obtain large virtual mass for the $q\bar{q}$ cluster comparable to the $Q_{0}$ of the probe. Analysis of the response function of the residual system indicates that the dominating mechanism is similar to the Feynman mechanism in which interacting parton carries almost all the momentum of the valence quark system, with the remaining light-front momentum distributed to the non-interacting wee partons.
The fit allows to describe the empirical PDF reasonably well in the range of $0.1\leq x\leq 0.85$ leaving little room for the hard mechanism of generating PDFs at large x. The model practically resolves the observed $(1-x)^{\beta}$, $\beta\approx 1$ scaling relation of the pion PDFs at large x.
With the parameters of $q\bar{q}$ and residual system Light-Front wave functions fixed one will be able to make numerical predictions for different observables such as generalized parton distribution functions (GPDs) or transverse momentum dependent parton distribution functions (TMDs) in pion that will allow for additional verification of the model.
Acknowledgement
This work is supported by United States Department of Energy grant under contract DE-FG02-01ER41172. AS acknowledges support through a Bridge to Doctorate NSF fellowship at FIU.
Appendix A Mandelstam s with many particles
For $n$ particles each with 4-momentum
$k^{\mu}_{i}=\left(k^{+}_{i},\frac{k_{\perp,i}^{2}+m^{2}_{i}}{k_{i}^{+}},\mathbf{k}^{\perp}_{i}\right)$ (46)
and total 4-momentum $k^{\mu}_{T}=\sum_{i=1}^{n}k^{\mu}_{i}=(x_{T}P^{+},k^{-}_{T},\mathbf{k}^{\perp}_{T})$, then the generalized Mandelstam $s_{n}$ is:
$s_{n}=\Bigg{(}\sum_{i}^{n}k^{\mu}_{i}\Bigg{)}^{2}=\sum_{i}^{n}\frac{\tilde{k}_{i,\perp}^{2}+m_{i}^{2}}{\beta_{i}}$ (47)
where:
$\beta_{i}=\frac{x_{i}}{x_{T}}$ (48)
$\mathbf{\tilde{k}}_{i,\perp}=\mathbf{k}_{i,\perp}-\beta_{i}\mathbf{k}_{T,\perp}$ (49)
The is seen by:
$s_{n}$ $=\Bigg{(}\sum_{i}^{n}k^{\mu}_{i}\Bigg{)}^{2}=\Bigg{(}\sum_{i}^{n}k^{+}_{i}\Bigg{)}\Bigg{(}\sum_{i}^{n}k^{-}_{i}\Bigg{)}-\Bigg{(}\sum_{i}^{n}\mathbf{k}_{i,\perp}\Bigg{)}^{2}$ (50)
$=(x_{T}P^{+})\Bigg{(}\sum_{i}^{n}\frac{{k}_{i,\perp}^{2}+m_{i}^{2}}{x_{i}P^{+}}\Bigg{)}-k^{2}_{T,\perp}$ (51)
$=\sum_{i}\frac{k^{2}_{i,\perp}+m^{2}_{i}}{\frac{x_{i}}{x_{T}}}-k^{2}_{T,\perp}$ (52)
Using the definition of $\beta_{i}$ and $\tilde{\mathbf{k}}_{i,\perp}$:
$s_{n}$ $=\sum_{i}\frac{\tilde{k}^{2}_{i,\perp}+2\beta_{i}\mathbf{k}_{T,\perp}\cdot\tilde{\mathbf{k}}_{i,\perp}+\beta_{i}^{2}k_{T,\perp}+m^{2}_{i}}{\beta_{i}}-k^{2}_{T,\perp}$ (53)
$=\sum_{i}\frac{\tilde{k}^{2}_{i,\perp}+2\beta_{i}\mathbf{k}_{T,\perp}\cdot\tilde{\mathbf{k}}_{i,\perp}+\beta_{i}^{2}k_{T,\perp}+m^{2}_{i}}{\beta_{i}}-k^{2}_{T,\perp}$ (54)
$=\sum_{i}\Bigg{(}\frac{\tilde{k}^{2}_{i,\perp}+m^{2}_{i}}{\beta_{i}}+2k_{T,\perp}\cdot\tilde{k}_{i,\perp}+\beta_{i}k^{2}_{T,\perp}\Bigg{)}-k^{2}_{T,\perp}$ (55)
Using the fact that $\sum_{i}\beta_{i}=1$ and $\sum_{i}\tilde{\mathbf{k}}_{i,\perp}=\sum_{i}({\mathbf{k}}_{i,\perp}-\beta_{i}\mathbf{k}_{T})=\mathbf{0}$,
$s_{n}=\sum_{i}\Bigg{(}\frac{\tilde{k}^{2}_{i,\perp}+m^{2}_{i}}{\beta_{i}}+k^{2}_{T,\perp}-k^{2}_{T,\perp}\Bigg{)}$ (56)
Then,
$s_{n}=\sum_{i}\frac{\tilde{k}^{2}_{i,\perp}+m^{2}_{i}}{\beta_{i}}$ (57)
Appendix B Light Front Energy Denominators
The LF energy denominator, $D$, for an intermediate state is given by:
$D$ $=\sum(k^{-}_{initial}-k^{-}_{intermediate})$ (58)
where $+$ and transverse momenta are conserved and we treat each particle on-mass shell, $k^{-}=\frac{m^{2}+k^{2}_{\perp}}{k^{+}}$. The denominator for nucleon splitting to the residual and valence parts is $D_{1}$:
$D_{1}$ $=p_{N}^{-}-k_{R}^{-}-k_{V}^{-}$
$=\frac{M^{2}_{N}}{p_{N}^{+}}-\frac{k_{R,\perp}+m^{2}_{R}}{k^{+}_{R}}-\frac{k_{V,\perp}+m^{2}_{V}}{k^{+}_{V}}$
$=\frac{1}{p^{+}_{N}}\left(M_{N}^{2}-\frac{k_{R,\perp}+m^{2}_{R}}{x_{R}}-\frac{k_{V,\perp}+m^{2}_{V}}{x_{V}}\right)$ (59)
Where $x_{R}=\frac{k^{+}_{R}}{p_{N}^{+}}$ and $x_{V}=\frac{k^{+}_{V}}{p_{N}^{+}}.$ The denominator for valence splitting to the 2 valence quarks of the Pion is $D_{2}$:
$D_{2}$ $=k^{-}_{V}+k^{-}_{R}-(k^{-}_{1i}+k^{-}_{2i}+k^{-}_{R})$
$=\frac{m^{2}_{V}+k_{V,\perp}^{2}}{k^{+}_{V}}-\frac{m^{2}_{1}+k_{1i,\perp}^{2}}{k^{+}_{1i}}-\frac{m^{2}_{2}+k_{2i,\perp}^{2}}{k^{+}_{2i}}$
$=\frac{1}{k^{+}_{V}}\left(m_{V}^{2}+k_{V,\perp}^{2}-\frac{k^{2}_{1i,\perp}+m^{2}_{1}}{\beta_{1i}}-\frac{k^{2}_{2i,\perp}+m^{2}_{2}}{\beta_{2i}}\right)$ (60)
Where $\beta_{1i}=\frac{k^{+}_{1i}}{k^{+}_{V}}$ and, $\beta_{2i}=\frac{k^{+}_{2i}}{k^{+}_{V}}$The denominator of the quark - quark interaction is $D_{3}$:
$D_{3}$ $=k^{-}_{1i}+k^{-}_{2i}-(k^{-}_{1m}+k^{-}_{2})$
$=\frac{1}{p_{N}^{+}}\Bigg{(}\frac{m^{2}_{1}+k_{1i,\perp}^{2}}{x_{1i}}+\frac{m^{2}_{2}+k_{2i,\perp}^{2}}{x_{2i}}$
$-\frac{m^{2}_{1}+k_{1m,\perp}^{2}}{x_{1m}}-\frac{m^{2}_{2}+k_{2,\perp}^{2}}{x_{2}}\Bigg{)}$
$=\frac{1}{p_{N}^{+}}\tilde{D_{3}}$ (61)